Jones Company has the following standards for its single product:

                     standard quantity            standard price
direct materials     11 pounds per unit           $4.25 per pound
direct labor          8 hours per unit            $14.00 per hour
variable overhead     8 hours per unit            ?????? per hour

Jones Company reported the following information for the month of October:

1.  9,140 units were produced.
2.  The direct material quantity variance was $36,295 favorable.
3.  The variable overhead spending variance was $1,520 favorable.
4.  The total direct labor variance was $7,140 unfavorable.
5.  Direct materials were purchased at a price of $4.60 per pound.
6.  The actual rate paid to direct laborers was $13.93 per hour.
7.  The actual variable overhead cost amounted to $627,480.
8.  At October 1, the direct materials inventory consisted of 2,000 pounds
    while the direct materials inventory at October 31 totaled 8,000 pounds.

Calculate the direct labor efficiency variance for October. If the variance
is favorable, place a minus sign in front of your answer (i.e., -5000). If
the variance is unfavorable, enter your answer as a number (i.e., 5000).
Answer #1

Standard direct labor cost = 9140*8*14 = $1023680

Actual direct labor cost = 1023680+7140 = 1030820

Actual hours = 1030820/13.93 = 74000 hours

Direct labor efficiency variance = (9140*8-74000)*14 = $12320 Unfavorable